need help solving this type of question
1/3:2/3=1/6:x

Answer by checkley77(12844)

1/3:2/3=1/6:x
1/3:2/3=1/6:x
x/3=2/3*1/6
x/3=2/18
18x=6
x=6/18
x=1/3 ans
Proof:
1/3:2/3=1/6:1/3
1/3*1/3=2/3*1/6
1/9=2/18
1/9=1/9